### v1

# CURATRON: Complete Robust Preference Data for Robust Alignment of Large Language Models

Abstract
This paper addresses the challenges of aligning large language models (LLMs) with human values via preference learning (PL), with a focus on the issues of incomplete and corrupted data in preference datasets. We propose a novel method for robustly and completely recalibrating values within these datasets to enhance LLMs resilience against the issues. In particular, we devise a guaranteed polynomial time ranking algorithm that robustifies several existing models, such as the classic Bradley–Terry–Luce (BTL) (Bradley and Terry,, 1952) model and certain generalizations of it. To the best of our knowledge, our present work is the first to propose an algorithm that provably recovers an $\epsilon$-optimal ranking with high probability while allowing as large as $O(n)$ perturbed pairwise comparison results per model response. Furthermore, we show robust recovery results in the partially observed setting. Our experiments confirm that our algorithms handle adversarial noise and unobserved comparisons well in both general and LLM preference dataset settings. This work contributes to the development and scaling of more reliable and ethically aligned AI models by equipping the dataset curation pipeline with the ability to handle missing and maliciously manipulated inputs.
[…]
Human feedback for open source LLMs needs to be crowd-sourced, Wikipedia style.
It is the only way for LLMs to become the repository of all human knowledge and cultures.
Yann LeCun
1 Introduction
Large Language Models (LLMs) are highly advanced Artificial Intelligence (AI) systems capable of understanding, interpreting, and generating languages. The integration of AI chatbots like ChatGPT into our daily lives and businesses has had a profound impact on both society and industries (Eloundou et al.,, 2023). These models have evolved from being specialized tools in specific fields to versatile assets that are increasingly applied in everyday activities and diverse work environments (Nguyen et al.,, 2023). However, the success of GPTs/LLMs depends not only on their ability to generate responses and perform tasks well but also on their alignment with human values and expectations.
Prior to ChatGPT, AI language models and chatbots often produced undesirable answers, making them unfit for public use. Examples of this problem include Microsoft’s Tay chatbot (Lee,, 2016), the National Eating Disorders Association’s Tessa (Mccarthy,, 2023), and Google’s Meena/LaMDA (Adiwardana et al.,, 2020). Although Meena/Lambda’s release was initially stopped due to not complying with Google’s AI safety and fairness guidelines, the project was later released publicly as Bard because of pressure from ChatGPT and Bing Chat (Mok,, 2023). Unfortunately, its factual error ended up costing Alphabet \$100 billion in market cap (Wittenstein,, 2023). The most recent AI incident at the time of writing this paper involved Gemini’s image generation AI model, which was found to be biased in favor of certain minority groups, leading Google to shut it down (Gordon,, 2024). Attempts have been made to address these problems by measuring the personality traits of large language models and modifying them through fine-tuning techniques (Karra et al.,, 2023). However, it was not until Instruct-GPT (Ouyang et al.,, 2022) was released with Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al.,, 2022), that marked a significant breakthrough in aligning LLMs effectively with human values. RLHF has paved the way for OpenAI’s ChatGPT to reshape industries, and GPT-4 to be considered a preliminary form of artificial general intelligence (AGI) (Bubeck et al.,, 2023), a type of AI that is vastly more intelligent than humans.
We show from the examples above that alignment is a crucial step in steering AI models toward the desired outcomes that align with human values/intentions. The prevalent method for aligning AI/LLMs currently involves preference learning (PL) through RLHF or Reinforcement Learning from AI Feedback (RLAIF) (Lee et al.,, 2023) using Proximal Policy Optimization (PPO) (Schulman et al.,, 2017), or alternatively, employing Direct Preference Optimization (DPO) (Rafailov et al.,, 2023). While PPO is a reinforcement learning (RL) technique within the RLHF pipeline, DPO directly integrates human preferences into the LLMs.
PL is an effective method, but it comes with a cost. This method requires high-quality human preference data (Touvron et al., 2023b, ). Acquiring human preference for LLMs is a time-consuming and expensive process that requires lots of resources and hiring qualified human labelers to ensure data quality (Bai et al.,, 2022; Lee et al.,, 2023). For LLaMa-2, while there was very limited information in the development of the model, the training cost was estimated to be between \$20-45 million, with \$8 million spent solely on preference data (Lambert,, 2023). Even for large AI companies like Google, Meta, OpenAI, and Anthtropic, it is often not feasible to account for all human values and cultures. Recent research by OpenAI shows that RLHF may not be able to scale to supervise AGI models reliably (Burns et al.,, 2023). One solution for the scaling problem is to crowdsource preference data collection to the general public, which significantly reduces costs and provides a more diverse set of human values. However, scaling up the collection of human preferences using this type of crowdsourcing often results in poor-quality data that is incomplete or adversarially corrupted. Participants may unintentionally or intentionally provide incomplete, inaccurate, or harmful feedback, as highlighted by (Casper et al.,, 2023; Draws et al.,, 2021; Daniel et al.,, 2018; Cocos et al.,, 2017). Using such data directly to train AI can result in suboptimal AI systems that are biased, have misaligned behaviors, or can lead to outcomes that are unfair, discriminatory, or harmful (Figure 1). Deploying models trained with poor-quality preference data can erode trust in AI systems, especially if they are used in critical applications. Users may lose confidence in the model’s ability to perform tasks accurately and align with their expectations and ethical standards, as mentioned in the examples above. Therefore, ensuring the reliability and integrity of preference datasets for aligning strong LLMs and AGI becomes challenging as the process scales up with large-scale responses and unscreened participants.
Approaching the issues, we consider the following learning problem. Suppose there are $n$ responses we wish to order based on a notion of comparison, between every pair of responses, with probabilistic outcomes. Further, we are given a set, =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$, consisting of $K$ independent pairwise comparison outcomes, denoted by $\{y_{ij}^{k}\}\in\{0,1\},k\in[K]$, between pairs of responses $(i,j)\subseteq[n]\times[n]$, a significant proportion of which might be corrupted by an adversary. In this passive learning setting, the concrete questions we wish to address are:
  1. 1.
    Is it possible to identify the pairs whose comparison results were corrupted by an adversary?
  2. 2.
    Having identified the corrupted results, as desired, is it possible to filter them out while computing a global ranking of the $n$ responses?
  3. 3.
    Is this task tractable statistically and computationally?
  4. 4.
    If so, is it possible to construct a provably correct and efficient algorithm, and what are the associated properties?
  5. 5.
    Further, does it work well in practice when we may also encounter unobserved data?
Our Contributions: We systematically answer the above questions in the affirmative. Specifically, our contributions are as follows.
Figure 1: CURATRON corrects incomplete and adversarially corrupted preference data to improve RLHF/DPO alignment results compared to using the raw initial preference data.
  1. 1.
    Problem formulation: We give a generic definition of (additive) adversarial noise, which can be handled for a broad class of statistical models, including the classic BTL model and also certain extensions of it such as the general Low–Rank (LR) models (Rajkumar and Agarwal,, 2016). As is the case with standard estimation techniques, if the noise is not modeled and handled well, we show that the quality of the estimated ranking could be quite bad, by quantifying the error of the estimated ranking with respect to the best possible ranking.
  2. 2.
    Algorithms & guarantees:
    • Under certain (information-theoretically tight) identifiability assumptions on the properties of the adversary, we develop a correct and efficient ranking method, Robust Preference Data for Reliably Alignment (RORATRON), that guarantees $\epsilon$-accurate high-probability learnability in a manner that is ‘robust’ and oblivious to the effects of the adversary. Our learning algorithm is provably characterized by polynomial time computational complexity.
    • In practice, it is often the case that not all pairs are compared, and even the observed pairwise comparison data could be adversarially corrupted – we also develop Complete Robust Preference Data for Reliable Alignment (CURATRON) and characterize the conditions for guaranteed robust recovery in this scenario. This results in a practical implication of enhancing preference data collection efficiency by automatically generating complete datasets from limited missing preference data.
  3. 3.
    Experiments: Finally, we support our theoretical results by showing robust ranking results on both synthetic and real-world experiments. Our experiments demonstrate the potential of our method in helping create large-scale AI/LLMs that are more accurately aligned with human values using minimal human effort as we achieved high reconstruction accuracy despite severe data missing and corruption.
2 Related Work
We now briefly present relevant work in: (1) LLM alignment with PL from human feedback, (2) ranking models and ranking algorithms that handle noise, and (3) robust subspace recovery methods, which will be needed for us to prove recovery results for ranking.
LLM Alignment with PL:
PL was initially developed to train agents in simulated environments to perform nuanced behaviors that are hard-to-define but easy to observe and recognize (Christiano et al.,, 2017). It has recently been found successful in aligning LLMs to human intentions and values such as harmfulness, helpfulness, factuality, and safety. Some of the methods of PL in LLMs are RLHF (Ouyang et al.,, 2022), DPO (Rafailov et al.,, 2023), Sequence Likelihood Calibration with Human Feedback (SLiC-HF) (Zhao et al.,, 2023), Rejection Sampling Optimization (RSO) (Liu et al., 2024b, ), and Pairwise Cringe Optimization (PCO) and iterative DPO (iDPO) (Xu et al.,, 2023). Recent works also attempt to eliminate the need for pairwise human preferences. Kahneman-Tversky Optimization (KTO) (Ethayarajh et al.,, 2023) requires only binary feedback on LLM outputs. Listwise Preference Optimization (LiPO) (Liu et al., 2024a, ) treats model alignment as a listwise ranking problem.
However, these methods require high-quality human supervision through pairwise or ranking preference data, which is often noisy and not practical in real-world scenarios (Casper et al.,, 2023). Trying to address this issue, robust DPO (rDPO) (Chowdhury et al.,, 2024), conservative DPO (cDPO) (Mitchell,, 2023), and Identity Preference Optimization (IPO) (Azar et al.,, 2023) losses are proposed to solve the problem of noisy preference data. An alternative solution to reduce dependence on human data is to utilize Constitutional AI/AIF (Bai et al.,, 2022; Lee et al.,, 2023). This approach aims to minimize human inputs by using other LLMs/models (Jiang et al., 2023b, ), or by self-critiquing and improving its outputs (Chen et al.,, 2024; Yuan et al.,, 2024) based on user-defined principles.
Our proposed approach to scaling up PL is different from previous work. Instead of creating new loss functions or using other models/LLMs themselves to remove humans outside of the loop, we aim to make PL with human feedback easier to manage by explicitly solving the problems associated with it. We develop low-cost and computationally efficient algorithms that take into account the low-rank characteristic of pairwise human preference data, which helps us address issues such as incomplete and adversarial corruptions in human preference datasets.
Ranking Models:
In the BTL model, item $i$ has an associated score $w_{i}$; then, the probability that item $i$ is preferred over $j$ is given by $P_{ij}=e^{-w_{i}}/(e^{-w_{i}}+e^{-w_{j}})$ where $\mathbf{w}\in\mathbb{R}^{n}$ is the BTL parameter vector to be estimated from data; here, $\mathbf{P}\in\mathbb{R}^{n\times n}$ is called the ‘preference matrix’. A closely related model, in the non-active setting, is the recently proposed LR model (Rajkumar and Agarwal,, 2016) wherein a generic class of preference matrices is characterized to be those having low rank under transformations using certain functions; specifically, for BTL-like models, the logit function defined as $\psi(x)=\log\left(x/(1-x)\right)$ turns out to right choice as shown in their paper. However, while their model accounts for missing information, they do not consider the harder problem of handling adversarial noise. Several robust ranking heuristics have been proposed (for example, (Wang et al.,, 2012; Zhou et al.,, 2011)), but these approaches do not have theoretical guarantees associated with them. The Sync-Rank algorithm proposed in (Cucuringu,, 2016) handles noise models different from the one considered in the preset work and is based on spectral techniques. Another related work is (Rajkumar and Agarwal,, 2014) which proposes the so-called ‘Generalized Low-Noise’ (GLN) condition that $\forall i\neq j,P_{ij}>P_{ji}\implies\sum_{h=1}^{n}\alpha_{h}P_{hj}>\sum_{h=1}^{n}\alpha_{h}P_{hi}$ for $\mathbf{\alpha}\in\mathbb{R}^{n}$. When $\alpha_{h}=1,\forall h$, they analyze the sample complexity and show convergence properties of various popular ranking algorithms like:
  1. 1.
    Maximum Likelihood (ML): this entails solving $\arg\max_{\mathbf{w}}\sum_{i<j}(\widehat{P}_{ij}(w_{j}-w_{i})-\log(1+\exp(w_{j}-w_{i})))$ where $\mathbf{w}\in\mathbb{R}^{n}$ is the BTL parameter vector and $\widehat{P}_{ij}$ is the empirical preference matrix.
  2. 2.
    Rank Centrality (RC) (Negahban et al.,, 2012): here, one sorts items by their scores which are computed as the stationary distribution of an appropriately normalized empirical preference matrix; this approach has a known sample complexity guarantee of $O(n\log(n))$.
  3. 3.
    Borda Count (BC) (Jiang et al.,, 2011): this heuristic involves ranking an item according to the fraction of times it beats other items.
For the general case $\mathbf{\alpha}$ (which previous methods fail to handle), they also propose a noise-tolerant SVM-based method for rank aggregation. However, in the adversarial setting, we consider in this paper, GLN could be violated and hence requires a different algorithmic approach and analysis.
Robust Subspace Recovery:
It is well-known that Principal Component Analysis (PCA), a ubiquitous technique for subspace identification, is not robust to outliers; this may be attributed to the fact that PCA is an $L_{2}$ optimization problem due to which grossly corrupted data points may perturb and skew the eigenvectors spanning the maximum variance subspace of the data points significantly.
The Robust PCA (RPCA) problem (Netrapalli et al.,, 2014) addresses the following question: suppose we are given a data matrix $\mathbf{M}$ which is the sum of an unknown low-rank matrix $\mathbf{L}$ and an unknown sparse matrix $\mathbf{S}$, can we recover each of the component matrices? RPCA is a widely used technique due to its ability to handle low-rankness, which is commonly observed in various applications such as recommendation systems, visual analytics, and social networks (Vaswani et al.,, 2018). Recently, RPCA has also been applied to parameter-efficient fine-tuning in the LLM domain (Nikdan et al.,, 2024).
While several works (Yi et al.,, 2016; Hsu et al.,, 2011) analyze the robust decomposition problem, it is shown in (Netrapalli et al.,, 2014) that, under information-theoretically tight assumptions, a simple iterative algorithm based on non-convex alternating projections of appropriate residuals provably yields an $\epsilon$-accurate solution in $O(\log(1/\epsilon))$ iterations with an overall computational complexity of $O(n^{2}r^{2}\log(1/\epsilon))$ where $r$ is the rank of $\mathbf{L}$. We will use this result, in particular, to derive guarantees for our ranking problem.
3 Problem Setup and Solution Approach
3.1 Notation
We first define some notation. We denote the set of all permutations of $n$ LLM responses/items as $\mathcal{S}_{n}$. If not specifically defined, we use lower-case letters for scalars, upper-case letters for global constants, lower-case bold-face letters for vectors and upper-case bold-face letters for matrices; specifically, $\mathbf{P}$ denotes a preference matrix. Let $\mathcal{P}_{n}:=\{\mathbf{P}\in[0,1]^{n\times n}|P_{ij}+P_{ji}=1\}$ denote the set of all pairwise preference matrices over $n$ responses. Let the set of stochastic-transitive matrices be $\mathcal{P}_{n}^{ST}:=\{\mathbf{P}\in\mathcal{P}_{n}|P_{ij}>1/2,P_{jk}>1/2\implies P_{ik}>1/2\}$. Let the set preference matrices described by the BTL model be $\mathcal{P}_{n}^{BTL}:=\{\mathbf{P}\in\mathcal{P}_{n}|\exists\mathbf{w}\in\mathbb{R}^{n}\text{ s.t. }e^{-w_{i}}/(e^{-w_{i}}+e^{-w_{j}})\}$. Let $\psi:[0,1]\mapsto\mathbb{R}$ be a strictly increasing bijective $L$-Lipschitz function and define the class of low-rank preference matrices with respect to $\psi$ as $\mathcal{P}_{n}^{LR(\psi,r)}=\{\mathbf{P}\in\mathcal{P}_{n}|\text{rank}(\psi(\mathbf{P}))\leq r\}$ where $r\in[n]$; when we apply such a transformation to a matrix, it is applied entry-wise. In this paper, we take $\psi$ to be the logit function.
For any matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$, let the infinity norm be denoted by $\left\lVert\mathbf{M}\right\rVert_{\infty}=\max_{i,j}\left\lvert M_{ij}\right\rvert$, the Frobenius norm be denoted by $\left\lVert\mathbf{M}\right\rVert_{F}=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{n}M_{ij}^{2}}$, the spectral norm be denoted by $\left\lVert\mathbf{M}\right\rVert_{2}=\max_{\mathbf{x,y}\in\mathbb{R}^{n}}\mathbf{x}^{\top}\mathbf{My}$. Denoting the indicator function by $\mathbbm{1}$, define the zero norm of a matrix to be the maximum number of non-zero elements in any row/column, ie, $\left\lVert\mathbf{M}\right\rVert_{0}=\max(\max_{j}\sum_{i=1}^{n}\mathbbm{1}(M_{ij}\neq 0),\max_{i}\sum_{j=1}^{n}\mathbbm{1}(M_{ij}\neq 0))$. Let the Singular Value Decomposition (SVD) of a square matrix be given by $\mathbf{M}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{\top}$ where $\mathbf{U,V}\in\mathbb{R}^{n\times r}$ are orthonormal matrices (whose columns are singular vectors) and $\bm{\Sigma}\in\mathbb{R}^{r\times r}$ is the diagonal matrix of singular values. Now, $\mathbf{M}$ is said to be $\mu$-incoherent if $\max\left(\max_{i}\left\lVert\mathbf{e}_{i}^{\top}\mathbf{U}\right\rVert_{2},\max_{i}\left\lVert\mathbf{e}_{i}^{\top}\mathbf{V}\right\rVert_{2}\right)\leq\mu\sqrt{r/n}$ where $\mathbf{e}_{i}$ denotes the $i^{th}$ basis vector in $\mathbb{R}^{n}$. Also, let $\sigma_{\max}:=\max_{i}\Sigma_{ii}$ and $\sigma_{\min}:=\min_{i}\Sigma_{ii}$.
We define the distance between a permutation $\sigma\in\mathcal{S}_{n}$ and a preference matrix $\mathbf{P}\in\mathcal{P}_{n}$ as:
$\operatorname{dist}\left(\sigma,\mathbf{P}\right)$ $:=\begin{pmatrix}n\\ 2\end{pmatrix}^{-1}\sum_{i<j}\mathbbm{1}\left((P_{ij}>1/2)\wedge(\sigma(i)\succ\sigma(j))\right)$
$+\begin{pmatrix}n\\ 2\end{pmatrix}^{-1}\sum_{i<j}\mathbbm{1}\left((P_{ji}>1/2)\wedge(\sigma(j)\succ\sigma(i))\right)$
Note that the above loss function basically is the number of pairs on which the ordering with respect $\sigma$ and $\mathbf{P}$ differ divided by the number of ways to choose two out of $n$ responses. Finally, let $P_{\min}=\min_{i\neq j}P_{ij}$ and $\Delta=\min_{i\neq j}\left\lvert\psi(P_{ij})-\psi(1/2)\right\rvert$.
3.2 Characterization of the Adversary
The following (weak) assumption characterizes the properties of the adversary. We shall see in the next section that it is information-theoretically tight in order to guarantee recovery in the solution approach that we propose. Note that this is a deterministic assumption; in particular, we do not have any distributional assumptions regarding the locations, the signs, or the magnitudes of the corruptions, and hence is very general.
Assumption 1.
The (additive) adversarial noise which corrupts a $\mu$-incoherent preference matrix $\mathbf{P}\in\mathcal{P}_{n}^{LR(\psi,r)}$ is modeled by a skew-symmetric sparse matrix $\mathbf{S}$ so that the corrupted preference matrix $\mathbf{P^{c}}\in\mathcal{P}_{n}$ is given by $\mathbf{P^{c}}=\mathbf{P}+\mathbf{S}$. We assume the (deterministic) bounded degree condition that $\left\lVert\mathbf{S}\right\rVert_{0}\leq d<n$ such $d<n/512\mu^{2}r$ where $r\leq n$.
Procedure 1 RPCA: Robust Principal Component Analysis
0:  $\mathbf{M}=\mathbf{L^{*}}+\mathbf{S^{*}}$, rank $r$ of $\mathbf{L^{*}}$.
0:  $\widehat{\mathbf{L}},\widehat{\mathbf{S}}$.
1:  Solve the following optimization problem using Algorithm 1 of (Netrapalli et al.,, 2014):
$\{\widehat{\mathbf{L}},\widehat{\mathbf{S}}\}=$ $\arg\min_{\mathbf{L},\mathbf{S}}\left\lVert\mathbf{M}-\mathbf{L}-\mathbf{S}\right\rVert_{F}$
[…]
2:  return  $\widehat{\mathbf{L}},\widehat{\mathbf{S}}$.
Procedure 2 PR: ($\gamma$-approximate) Pairwise Ranking
0:  Preference matrix $\mathbf{M}\in\mathbb{R}^{n\times n}$.
0:  Ranking $\widehat{\sigma}$.
1:  Compute $\forall i,\quad v_{i}\leftarrow\sum_{j=1}^{n}\mathbbm{1}(M_{ij}>1/2)$.
2:  return  $\widehat{\sigma}\leftarrow\text{Sort}(\mathbf{v})$.
So, why do existing non-robust algorithms not recover the true response ordering in the presence of an adversarial noise source? This question is answered by the following proposition which precisely quantifies how bad a ranking could be when an algorithm uses the corrupted pairwise preference matrix. The key idea is to construct an adversary that intentionally flips true comparison results.
Claim 1 (Upper bound on estimation error).
Under Assumption 1 it is possible that $\operatorname{dist}(\widehat{\sigma},\mathbf{P^{c}})=O(1)$.
Proof.
Assume that we are exactly given the entries of the preference matrix as opposed to sampling them. Note that in order to estimate a ranking from a given preference matrix, we still need to use a pairwise ranking procedure. Let $\widehat{\sigma}\in\mathcal{S}_{n}$ be the output of any Pairwise Ranking (PR) procedure with respect to an underlying preference matrix $\mathbf{Q}\in\mathcal{P}_{n}$. For a constant $\gamma>1$, $\widehat{\sigma}$ is said to be $\gamma$-approximate if $\operatorname{dist}(\widehat{\sigma},\mathbf{Q})\leq\gamma\min_{\sigma\in\mathcal{S}_{n}}\operatorname{dist}(\sigma,\mathbf{Q})$. Define the following distance which measures the fraction of response pairs over which two preference matrices $\{\mathbf{Q},\mathbf{R}\}\in\mathcal{P}_{n}$ disagree.
$\operatorname{dist}\left(\mathbf{Q},\mathbf{R}\right)$ $:=\begin{pmatrix}n\\ 2\end{pmatrix}^{-1}\sum_{i<j}\mathbbm{1}\left((Q_{ij}>1/2)\wedge(R_{ij}<1/2)\right)$
$+\begin{pmatrix}n\\ 2\end{pmatrix}^{-1}\sum_{i<j}\mathbbm{1}\left((Q_{ij}<1/2)\wedge(R_{ij}>1/2)\right)$
By Lemma 20 of (Rajkumar and Agarwal,, 2016), for $\mathbf{Q}\in\mathcal{P}_{n}^{ST}$ and $\mathbf{R}\in\mathcal{P}_{n}$, we have $\operatorname{dist}(\widehat{\sigma},\mathbf{Q})\leq(1+\gamma)\operatorname{dist}(\mathbf{Q},\mathbf{R})$. But note that it is possible that $\operatorname{dist}\left(\mathbf{Q},\mathbf{R}\right)=1$ as it is easy to construct by $\mathbf{R}$ that disagrees with $\mathbf{Q}$ in every entry by simply setting $\mathbf{R}=\mathbf{Q}^{\top}$. Now, we may set $\mathbf{Q}=\mathbf{P}$ and $\mathbf{R}=\mathbf{P^{c}}$ for any algorithm that uses $\mathbf{P^{c}}$ for ranking; specifically, for the adversary satisfying Assumption 1, we can see by a direct counting argument that $\operatorname{dist}\left(\mathbf{Q},\mathbf{R}\right)\leq\frac{d(2n-1-d)}{n(n-1)}$ which proves the claim. ∎
3.3 Alignment of LLMs using Direct Preference Optimization
We utilize Direct Preference Optimization (DPO) (Rafailov et al.,, 2023) to align LLMs in our experiment to examine the efficacy of our algorithms in solving incomplete and adversarially corrupted preference data problems. DPO is a widely used method for PL, due to its simplicity, stability, performance, and computational efficiency. Essentially, DPO is a loss function that can be expressed as a mathematical equivalent to Reinforcement Learning with Human Feedback (RLHF). Its purpose is to optimize preferences by classifying human preference data without resorting to reward modeling or reinforcement learning.
$\mathcal{L}_{DPO}(\theta,y_{w},y_{l})=-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w})}{\pi_{\text{ref}}(y_{w})}-\beta\log\frac{\pi_{\theta}(y_{l})}{\pi_{\text{ref}}(y_{l})}\right),$
where $y_{w}$ and $y_{l}$ represent preferred and dispreferred responses respectively. $\pi_{\text{ref}}$ is the reference policy, and $\pi_{\theta}$ is the new policy.
3.4 Solution Approach
In real-world scenarios, it is common to have missing or corrupted data, especially in large-scale crowdsource data. Our proposed methods aim to take the raw corrupted data as input and output the recovered data that closely resembles the counterfactual original data. This recovered data is then input into the alignment step, as illustrated in Figure 1. To achieve this, we first transform the raw preference data into a logit preference matrix space to utilize the low-rank characteristic. Different ranking approaches, such as, pointwise, pairwise, and listwise, can be represented in this logit preference matrix space.
${{\psi}(P)=\log{\frac{P}{1-P}}},\hskip 5.0pt\text{where}\hskip 5.0ptP={\frac{e^{-w_{i}}}{e^{-w_{i}}+e^{-w_{j}}}}$
We now identify three scenarios/settings where missing and adversarially corrupted comparisons can affect the ranking of responses and tackle them in detail in subsequent sections.
  1. 1.
    Fully observed and adversarially corrupted setting (Figure 1(a), Section 4) : Some instances of the comparison results are adversarially corrupted. This scenario can happen when data quantity is prioritized over data quality in the data collection process, resulting in biased or malicious human feedback. Algorithm 3 RORATRON is proposed to solve this problem.
  2. 2.
    Partially observed and uncorrupted setting (Figure 1(b), Section 5.1) : Not all pairs of responses are compared. This scenario can happen when data quality is prioritized over quantity in the data collection process. Observing all possible comparisons can be expensive and challenging, especially when there are many LLM responses to compare. Algorithm 4 CORATRON is proposed to solve this problem.
  3. 3.
    Partially observed and adversarially corrupted setting (Figure 1(c), Section 5.2) : Both (1) not all pairs of responses are compared, and (2) some instances of the comparison results are adversarially corrupted. This scenario can happen when data quantity and quality are not met in the data collection process. This scenario will likely happen in a large crowd-sourced environment due to large-scale LLM responses and participants. Algorithm 5 CURATRON is proposed to solve this problem.
Figure 2: Three settings/scenarios where incomplete data and adversarial corruption can affect LLMs: (a) Fully observed and adversarially corrupted setting, (b) Partially observed and uncorrupted setting, and (c) Partially observed and adversarially corrupted setting.
4 Fully Observed and Adversarially Corrupted Setting
4.1 Algorithm
In this section, we answer Question 4. We present our main algorithm for robust passive ranking from pairwise comparisons in the presence of adversarial noise in Algorithm 3. The input data consist of the set of pairwise comparison results =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$, $(i,j)\in[n]\times[n]$, $k\in[K]$, $y_{ij}^{k}\in\{0,1\}$. The algorithm assumes the true rank of $\psi(\mathbf{P})$ as an input parameter; specifically, for the BTL model, we set $r=2$. Algorithm 3 RORATRON calls the following procedures:
  1. 1.
    Robust PCA (Procedure 1): Note that Step 3 of Algorithm 3 uses a matrix low-rank plus sparse decomposition subroutine. To obtain our recovery guarantee, it is sufficient to use the robust PCA problem as a black-box method; for the precise details of this algorithm, we refer the reader to (Netrapalli et al.,, 2014). In particular, for our analysis, we use the noise-case guarantees in their paper. This is charaterized by a (strongly-polynomial) running time of $O(n^{2}r^{2}\log(1/\epsilon))$ and guarantees $\epsilon$-recovery of the component matrices under the conditions of Assumption 1 and Lemma 3.
  2. 2.
    $\gamma$-approximate pairwise ranking procedure (Subroutine 2): Step 4 of Algorithm 3 calls a constant factor approximate ranking procedure. Specifically, we use the Copeland procedure (Copeland,, 1951) which has a $5$-approximation guarantee (Coppersmith et al.,, 2006) and involves sorting the responses according to a score of response $i$ given by $\sum_{j=1}^{n}\mathbbm{1}(\overline{P}_{ij}>1/2)$.
4.2 Analysis
Algorithm 3 RORATRON: Robust Preference Data for Robust Alignment
0:  Comparison dataset =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$, true rank $r$.
0:  Ranking of $n$ responses, $\hat{\sigma}\in\mathcal{S}_{n}$.
1:  Estimate entries of $\mathbf{\widehat{P}}$ for $i\leq j$ as:
$\widehat{P}_{ij}=\begin{cases}\frac{1}{K}\sum_{k=1}^{K}y_{ij}^{k}&\text{ if }i<j\\ 1/2&\text{ if }i=j\\ \end{cases}$
2:  Set $\widehat{P}_{ij}=1-\widehat{P}_{ji}$ for all $i>j$.
3:  Perform robust PCA: $\{\mathbf{\psi(\overline{P})},\mathbf{\widehat{S}}\}\leftarrow\text{RPCA}(\mathbf{\psi(\widehat{P})},r)$.
4:  Using a pairwise ranking procedure after taking the inverse transform: $\widehat{\sigma}\leftarrow\text{PR}(\mathbf{\overline{P}})$.
5:  return  $\widehat{\sigma}$.
We begin with a useful short result followed by the statement and the proof of our main result that, with high probability, we achieve $\epsilon$–accurate ranking in polynomial time using polynomial number of samples, despite the presence of adversarial noise. Precisely, Theorem 1 and Remark 1 address Question 3; Remark 2 addresses Question 1. In this context it is noteworthy that we present the result for LR models which strictly contain the BTL model while being much more general (Rajkumar and Agarwal,, 2016); upon proving this result, we specialize it to the classic BTL model as well (Corollary 1).
Lemma 1 (Some properties of the logit function).
Let $a,b,c\in(0,1)$ such that $c=a+b$. Then, we have,
  1. 1.
    $\psi(c)=\psi(a)+\psi(a+b)+\psi(1-a)$
  2. 2.
    $\psi(a)+\psi(1-a)=0$.
Proof.
Both follow by using the definition of the logit function that $\psi(a)=\log(a/(1-a))$ and using the property that $\log(ab)=\log(a)+\log(b)$. ∎
Theorem 1 (Provably good estimation of ranking in LR models in the presence of adversarial noise).
Let $\mathbf{P}\in\mathcal{P}_{n}^{LR(\psi,r)}$ be the true preference matrix according to which the pairwise comparison dataset =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$ is generated for all responses pairs $(i,j)$ such that $k\in[K]$. Let $\mathbf{\widehat{P}}$ be the empirical preference matrix computed using $\aleph$. Let $\mathbf{S}\in[0,1]^{n\times n}$ be the adversarial matrix that additively corrupts $\mathbf{\widehat{P}}$. Let $\psi$ be $L$-Lipschitz in $[\frac{P_{\min}}{2},1-\frac{P_{\min}}{2}]$ and $\psi(\mathbf{P})$ be $\mu$-incoherent. Let each pair be compared independently $K\geq 16384\mu^{2}(1+\gamma)L^{2}n^{2}\log^{2}(n)/\epsilon\Delta^{2}$ times where $\Delta=\min_{i\neq j}\left\lvert\psi(P_{ij})-\psi(1/2)\right\rvert$. Then, with probability at least $1-1/n^{3}$, Algorithm 3 returns an estimated permutation $\widehat{\sigma}$ such that $\operatorname{dist}(\widehat{\sigma},\mathbf{P})\leq\epsilon$.
Remark 1 (Computational complexity).
In Algorithm 3, Step 1 takes $O(n^{2}K)=O(n^{4}\log^{2}n/\epsilon)$ time, Step 3 takes $O(n^{2}r^{2}\log(1/\epsilon))$, and Step 4 takes $O(n^{2}+n\log n)$ time. Thus, putting together the cost of these main steps, the overall computational complexity of our robust ranking algorithm for $\mathbf{P}\in\mathcal{P}_{n}^{LR(\psi,r)}$ is $O(n^{4}\log^{2}n/\epsilon)$.
Remark 2 (Identifying adversarially corrupted pairwise comparisons).
From Step 3 of Algorithm 3, using Theorem 2 of (Netrapalli et al.,, 2014), we also have $\operatorname{Supp}(\mathbf{\widehat{S}})\subseteq\operatorname{Supp}(\mathbf{S})$ and thus we can identify the corrupted pairwise comparison results.
Remark 3 (Missing data versus adversarially corrupted data).
Note that the adversarial sparse noise we consider subsumes the setting when comparison results for certain pairs are missing as in (Rajkumar and Agarwal,, 2016) and hence directly applies in that situation. Moreover, since the support and magnitude of the corrupted entries of the preference matrix are unknown, the problem considered in this paper is harder; consequently, our sample complexity is $O(n^{2})$ as opposed to $O(n\operatorname{poly}\log n)$ in their work.
Proof.
Let $\widetilde{P}_{ij}$ be the empirical probability estimate of $P_{ij}$. Note that we compute $\widetilde{P}_{ij}=\frac{1}{K}\sum_{k=1}^{K}y_{ij}^{k}$ from the given pairwise comparison dataset, =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$. Now, $\mathbf{\widehat{P}}=\mathbf{\widetilde{P}}+\mathbf{S}$. By Lemma 1, we may write the adversarially corrupted empirical probability estimate as $\psi(\mathbf{\widehat{P}})=\psi(\mathbf{\widetilde{P}})+\mathbf{\widetilde{S}}$ where $\mathbf{\widetilde{S}}=\psi(\mathbf{\widetilde{P}}+\mathbf{S})+\psi(1-\mathbf{\widetilde{P}})$. We have $\psi(\mathbf{\widetilde{P}})=\psi(\mathbf{P})+\mathbf{\widetilde{N}}$ where $\mathbf{\widetilde{N}}=\psi(\mathbf{\widetilde{P}})-\psi(\mathbf{P})$. Now, this noise, $\mathbf{\widetilde{N}}$, is purely due to finite-sample effects which can be controlled (using concentration arguments given in the inequality $\xi_{3}$ below) by driving it down to as small a value as we want by ensuring large enough number of comparisons for each pair. Note that we input $\psi(\mathbf{\widehat{P}})=\psi(\mathbf{P})+\mathbf{\widetilde{S}}+\mathbf{\widetilde{N}}$ to Subroutine 1 and obtain $\psi(\mathbf{\overline{P}})$ as the output in Step 3 of Algorithm 3. Hence, using Theorem 2 from (Netrapalli et al.,, 2014), if $\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}\leq\sigma_{\min}(\psi(\mathbf{P}))/100n$, we have,
$\left\lVert\psi(\mathbf{\overline{P}})-\psi(\mathbf{P})\right\rVert_{F}\leq\epsilon^{\prime}+2\mu^{2}r\left(7\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}+\frac{8n}{r}\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}\right)$
after $T\geq 10\log(3\mu^{2}r\sigma_{1}/\epsilon^{\prime})$ iterations associated with Step 1 of Subroutine  1. Next, we have, with probability at least $1-1/n^{3}$,
$\left\lVert\psi(\mathbf{\overline{P}})-\psi(\mathbf{P})\right\rVert_{F}$ $\leq\epsilon^{\prime}+2\mu^{2}r\left(7\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}+\frac{8n}{r}\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}\right)$
$\stackrel{{\scriptstyle\xi_{1}}}{{\leq}}\epsilon^{\prime}+32\mu^{2}n\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}\stackrel{{\scriptstyle\xi_{2}}}{{\leq}}\epsilon^{\prime}+32\mu^{2}n\tau$
$\stackrel{{\scriptstyle\xi_{3}}}{{\leq}}n\sqrt{\frac{\epsilon}{1+\gamma}}\frac{\Delta}{2}$
where $\xi_{1}$ follows by using $r\leq n$ and $\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}\leq\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}$, $\xi_{2}$ follows by substituting for $\mathbf{\widetilde{N}}$ from Lemma 2 with $K\geq\frac{L^{2}n^{2}\log^{2}n}{\tau^{2}}$, and $\xi_{3}$ is obtained using $\epsilon^{\prime}=n\sqrt{\frac{\epsilon}{1+\gamma}}\frac{\Delta}{4}$, $\tau=\min\left(\sigma_{\min}(\psi(\mathbf{P}))/100,\sqrt{\frac{\epsilon}{1+\gamma}}\frac{\Delta}{128\mu^{2}}\right)$. Then using similar arguments as proof of Theorem 13 in (Rajkumar and Agarwal,, 2016), we obtain our result. ∎
Lemma 2 (Concentration of sampling noise).
Under the conditions of Theorem 1, let each response pair be compared such that the number of comparisons per response pair is $K\geq\frac{L^{2}n^{2}\log(n)}{\tau^{2}}$; with probability at least $1-1/n^{3}$, $\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}\leq\tau$.
Proof.
Let $L$ be the Lipschitz constant of $\psi$ and set $K\geq\frac{L^{2}n^{2}\log(n)}{\tau^{2}}$. Using the inequality that $\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}\leq n\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}$,
$\Pr\left(\left\lVert\mathbf{\widetilde{N}}\right\rVert_{2}\geq\tau\right)$ $\leq\Pr\left(\left\lVert\mathbf{\widetilde{N}}\right\rVert_{\infty}\geq\frac{\tau}{n}\right)$
$=\Pr\left(\exists(i,j):\left\lvert\psi(\widehat{P}_{ij})-\psi(P_{ij})\right\rvert\geq\frac{\tau}{n}\right)$
$\leq\sum_{i,j}\Pr\left(\left\lvert\psi(\widehat{P}_{ij})-\psi(P_{ij})\right\rvert\geq\frac{\tau}{n}\right)$
$\leq\sum_{i,j}\Pr\left(\left\lvert\widehat{P}_{ij}-P_{ij}\right\rvert\geq\frac{\tau}{nL}\right)\leq\frac{1}{n^{3}}$
Next, for completeness, we recall the following lemma (proved in Theorem 8 and Lemma 14 of (Rajkumar and Agarwal,, 2016)) which characterizes the incoherence constant $\mu$ of $\mathbf{P}\in(\mathcal{P}_{n}^{LR(\psi,2)}\cap\mathcal{P}_{n}^{ST})$ in Assumption 1.
Lemma 3 (Incoherence of BTL and LR models).
We have $\mathbf{P}\in(\mathcal{P}_{n}^{LR(\psi,2)}\cap\mathcal{P}_{n}^{ST})$ if and only if $\psi(\mathbf{P})=\mathbf{uv}^{\top}-\mathbf{vu}^{\top}$ for $\mathbf{u}\in\mathbb{R}^{n}_{+}$ and $\mathbf{v}\in\mathbb{R}^{n}$ where $\mathbf{u}^{\top}\mathbf{v}=0$. Moreover, $\psi(\mathbf{P})$ is $\mu$-incoherent where $\mu=\sqrt{\frac{n}{2}}\left(\frac{u_{\max}^{2}}{u_{\min}^{2}}+\frac{v_{\max}^{2}}{v_{\max}^{2}}\right)^{1/2}$ where $u_{\min}=\min_{i}\left\lvert u_{i}\right\rvert$, $u_{\max}=\max_{i}\left\lvert u_{i}\right\rvert$, $v_{\min}=\min_{i}\left\lvert v_{i}\right\rvert$ and $v_{\max}=\min_{i}\left\lvert v_{i}\right\rvert$. We also have $\mathcal{P}_{n}^{BTL}\subset(\mathcal{P}_{n}^{LR(\psi,2)}\cap\mathcal{P}_{n}^{ST})$ since we may set $\mathbf{u}=\mathbf{1}$ where $\mathbf{1}$ is the all-ones vector and $\mathbf{v}=\mathbf{w}$ where $\mathbf{w}$ is the BTL parameter vector. In this case, we may rewrite $\mu=\sqrt{\frac{n}{2}}\left(1+\frac{(w_{\max}-\overline{w})^{2}}{(w_{\min}-\overline{w})^{2}}\right)$ where $\overline{w}=\frac{1}{n}\sum_{i=1}^{n}w_{i}$.
The following corollary makes precise our claim that up to $O(n^{2})$ response pairs may be subject to adversarial corruption but our RORATRON algorithm still recovers a good ranking.
Corollary 1 (Recovery result for BTL model).
Consider $\mathbf{P}\in\mathcal{P}_{n}^{BTL}$. Using Assumption 1, let the adversarial matrix be $\mathbf{S}\in[0,1]^{n\times n}$ satisfying $\left\lVert\mathbf{S}\right\rVert_{0}\leq n/1024\mu^{2}$ where $\mu$ is characterized as in Lemma 3. Then, with probability $1-1/n^{3}$, the output of Algorithm 3 with input $\mathbf{\widehat{P}}$ computed using =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$ satisfies and $r=2$, $\operatorname{dist}(\widehat{\sigma},\mathbf{P})\leq\epsilon$.
5 Partially Observed Uncorrupted and Adversarially Corrupted Settings
5.1 Partially Observed and Uncorrupted Setting
In the case when we are confident that the data are collected faithfully but we do not have the complete data to work with, we can use the ‘OptSpace’ matrix completion algorithm of (Keshavan et al.,, 2010) by itself to generate the full preference matrix from the incomplete one, as presented in Algorithm 4 CORATRON. We show in Experiment 7.3 below that we can still complete the full matrix with minimal error in such a setting with extremely missing data.
Algorithm 4 CORATRON: Complete Preference Data for Robust Alignment
0:  Comparison dataset =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$, true rank $r$.
0:  Ranking of $n$ responses, $\hat{\sigma}\in\mathcal{S}_{n}$.
1:  Estimate entries of $\mathbf{\widehat{P}}$ for $i\leq j$ as:
$\widehat{P}_{ij}=\begin{cases}\frac{1}{K}\sum_{k=1}^{K}y_{ij}^{k}&\text{ if }i<j\text{ and }(i,j)\in\Omega\\ 1/2&\text{ if }i=j\text{ and }(i,j)\in\Omega\\ 1/2&\text{ if }(i,j)\notin\Omega\\ \end{cases}$
2:  Set $\widehat{P}_{ij}=1-\widehat{P}_{ji}$ for all $i>j$.
3:  Set $\mathbf{R}\leftarrow\text{OptSpace}(\psi(\mathbf{\widehat{P}})_{\Omega})$.
4:  Using a pairwise ranking procedure after taking the inverse transform: $\widehat{\sigma}\leftarrow\text{PR}(\mathbf{R})$.
5:  return  $\widehat{\sigma}$.
5.2 Partially Observed and Adversarially Corrupted Setting
In this subsection, we consider the partially observed and adversarially corrupted comparison results setting. Both factors can be modeled in a unified manner by setting the corresponding missing entries of the preference matrix to zero (or a specific constant to account for numerical stability). We present our robust ranking algorithm for this setting in Algorithm 5 CURATRON – this essentially involves using OptSpace (Keshavan et al.,, 2010) followed by using the robust PCA algorithm of (Netrapalli et al.,, 2014) as sub-routines. We note at this point that, while (Niranjan and Rajkumar,, 2017) work considers the incomplete data case, it leverages extra information provided in the form of side information (specifically, noiseless and complete item-related features) to derive recovery guarantees. Also, their algorithm is still unable to handle the presence of pairwise comparisons corrupted in an adversarial manner, which is explored in the (Niranjan et al.,, 2017) work. We now derive the recovery guarantees as follows.
Algorithm 5 CURATRON: Complete Robust Preference Data for Robust Alignment
0:  Comparison dataset =$\aleph=\{(i,j,\{y_{ij}^{k}\})\}$, true rank $r$.
0:  Ranking of $n$ responses, $\hat{\sigma}\in\mathcal{S}_{n}$.
1:  Estimate entries of $\mathbf{\widehat{P}}$ for $i\leq j$ as:
$\widehat{P}_{ij}=\begin{cases}\frac{1}{K}\sum_{k=1}^{K}y_{ij}^{k}&\text{ if }i<j\text{ and }(i,j)\in\Omega\\ 1/2&\text{ if }i=j\text{ and }(i,j)\in\Omega\\ 1/2&\text{ if }(i,j)\notin\Omega\\ \end{cases}$
2:  Set $\widehat{P}_{ij}=1-\widehat{P}_{ji}$ for all $i>j$.
3:  Set $\mathbf{R}\leftarrow\text{OptSpace}(\psi(\mathbf{\widehat{P}})_{\Omega})$.
4:  Use a robust PCA procedure: $\mathbf{\psi(\overline{P})}\leftarrow\text{RPCA}(\mathbf{R})$.
5:  Using a pairwise ranking procedure after taking the inverse transform: $\widehat{\sigma}\leftarrow\text{PR}(\mathbf{\overline{P}})$.
6:  return  $\widehat{\sigma}$.
Theorem 2 (Provably good estimation of ranking in BTL model in the presence of adversarial noise as well as missing data).
Consider similar notations as in Theorem 1 but let $\mathbf{P}\in\mathcal{P}_{n}^{BTL}$. Let $\Omega\subseteq[n]\times[n]$ be a set of compared response pairs. Assume $\Omega$ is drawn uniformly from all subsets of $[n]\times[n]$ of size $\left\lvert\Omega\right\rvert$ such that $\left\lvert\Omega\right\rvert\geq C^{\prime\prime}n\log(n)$ and let the sparse noise satisfy $\left\lVert\mathbf{S}\right\rVert_{\infty}\leq\Delta_{w}\frac{\log(n)}{C_{\Delta}n}$ where $\Delta_{w}:=\min_{i,j}\left\lvert w_{i}-w_{j}\right\rvert$. Let the number of comparisons per pair be $K\geq cn^{4}/\Delta_{w}$. Then with probability at least $1-2/n^{3}$, Algorithm 5 returns a ranking that satisfies $\operatorname{dist}(\widehat{\sigma},\mathbf{P})\leq\epsilon$.
Remark 4 (Robust Estimation of BTL Model in the Partially Observed Case).
For the BTL model, Theorem 2 says $O(n\log n)$ pairs suffice to estimate the BTL model, which matches bounds from (Rajkumar and Agarwal,, 2016). Further, even in this incomplete comparison data case, we are able to tolerate uniformly random additive sparse noise with its maximum absolute entry scaling as the order of the BTL ‘score-gap’ divided by the number of responses up to logarithmic factors, ie, $\widetilde{O}\left(\Delta_{w}/n\right)$.
Proof.
From Lemma 3, we have $\psi(\mathbf{P})=\mathbf{1w}^{\top}-\mathbf{w1}^{\top}$ for the BTL model where $\psi$ is the logit function. Clearly, in this case, $\psi(\mathbf{P})$ is a real skew-symmetric matrix of rank $r=2$. Since it is skew-symmetric, its eigenvalues, which are the roots of its characteristic polynomial, are of the form ±$\pm\lambda i$ for some $\lambda\in\mathbb{R}$ and $i=\sqrt{-1}$, and hence, $\sigma_{\min}(\psi(\mathbf{P}))=\sigma_{\max}(\psi(\mathbf{P}))$, ie, the condition number of $\psi(\mathbf{P})$, $\kappa=1$. Now, we recall the spectral-lower bound from Corollary 2 of (Horne,, 1997),
$\sigma_{\min}(\psi(\mathbf{P}))\geq\frac{\left\lVert\psi(\mathbf{P})\right\rVert_{F}}{\sqrt{r(r-1)}}\geq\sqrt{\frac{n(n-1)}{2}}\Delta_{w}$ (1)
where $\Delta_{w}=\min_{i,j}\left\lvert w_{i}-w_{j}\right\rvert$.
Let $\Omega\subseteq[n]\times[n]$ be a subset of all the response pairs with comparison results among which some might be corrupted by sparse noise, ie, $\psi(\mathbf{\widehat{P}}_{\Omega})=\psi(\mathbf{P}_{\Omega})+\mathbf{\widetilde{S}}_{\Omega}+\mathbf{\widetilde{N}}_{\Omega}$. Let $\mathbf{T}:=\mathbf{\widetilde{S}}_{\Omega}+\mathbf{\widetilde{N}}_{\Omega}$. From Theorem 1.2 of (Keshavan et al.,, 2010), we have $\frac{1}{n}\left\lVert\psi(\mathbf{\widehat{P}})-\psi(\mathbf{P})\right\rVert_{F}=\frac{1}{n}\left\lVert\mathbf{T}+\mathbf{M}\right\rVert_{F}\leq C\kappa^{2}\frac{n\sqrt{r}}{\left\lvert\Omega\right\rvert}\left\lVert\mathbf{T}\right\rVert_{2}$ where $\mathbf{M}$ is the noise matrix after obtaining the completed matrix $\psi(\mathbf{\widehat{P}})$ from $\psi(\mathbf{\widehat{P}}_{\Omega})$ using OptSpace. Using triangle inequality and noting that $\left\lvert\Omega\right\rvert\geq C^{\prime\prime}n\log(n)$, the noise may be bounded as
$\left\lVert\mathbf{\widetilde{N}}_{\Omega}+\mathbf{M}\right\rVert_{\infty}$ $\leq\left\lVert\mathbf{\widetilde{N}}_{\Omega}+\mathbf{M}\right\rVert_{F}\leq\left\lVert\mathbf{T}\right\rVert_{2}\frac{\sqrt{2}Cn^{2}}{\left\lvert\Omega\right\rvert}+\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{F}$
$\stackrel{{\scriptstyle\zeta_{1}}}{{\leq}}C^{\prime}\frac{n}{\log(n)}\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{2}$ (2)
where $C$, $C^{\prime}$ and $C^{\prime\prime}$ are constants and $\zeta_{1}$ is obtained by using the triangle inequality that $\left\lVert\mathbf{T}\right\rVert_{2}\leq\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{2}+\left\lVert\mathbf{\widetilde{N}}_{\Omega}\right\rVert_{2}$, followed by setting $K\geq cn^{4}/\Delta_{w}$ for constant $c$ and finally using $\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{F}\leq\sqrt{n}\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{2}$. Then, combining Equations 2 and 1, we have if
$\frac{\log(n)}{C_{\Delta}n}\Delta_{w}$ $\geq\left\lVert\mathbf{\widetilde{S}}_{\Omega}\right\rVert_{2}=\left\lVert\psi(\mathbf{\widehat{P}})-\psi(\mathbf{\widetilde{P}})\right\rVert_{2}$
$\geq\left\lVert\psi(\mathbf{\widehat{P}})-\psi(\mathbf{\widetilde{P}})\right\rVert_{\infty}\geq L\left\lVert\mathbf{\widehat{P}}-\mathbf{\widetilde{P}}\right\rVert_{\infty}\geq\left\lVert\mathbf{S}\right\rVert_{\infty}$
where $C_{\Delta}$ is a global constant and using Lemma 2, then we have the guarantee (along similar lines as that of Theorem 1 that Algorithm 5 returns an estimated permutation which satisfies $\operatorname{dist}(\widehat{\sigma},\mathbf{P})\leq\epsilon$. ∎
6 Generalization to Other Ranking Models
Related to the BTL model are many other binary choice models (Fishburn,, 1973) such as the Thurstonian model (Thurstone,, 1927). In such models, the preference matrix has been shown to be low-rank under appropriate choices of $\psi$; for instance, for the Thurstonian models, the probit function turns out to be the right choice. For further details, we refer the reader to the work of (Rajkumar and Agarwal,, 2016).
Let $a,b,c\in(0,1)$ such that $c=a+b$. Then, for any general non-linear $L$-Lipschitz function, we write $\psi(c)=\psi(a+b)=\psi(a)+\psi(a+b)-\psi(a)$. The error may be lower bounded by $\left\lvert\psi(a+b)-\psi(a)\right\rvert\geq Lb$. Thus, for any adversarial model wherein we have $\mathbf{P^{c}}=\mathbf{P}+\mathbf{S}$, we have:
$\psi(\mathbf{P^{c}})$ $=\psi(\mathbf{P})+(\psi(\mathbf{P}+\mathbf{S})-\psi(\mathbf{P}))=\psi(\mathbf{P})+\mathbf{\widetilde{S}}$
where $\mathbf{\widetilde{S}}$ is also a deterministic sparse corruption matrix with the absolute value of the non-zero entries lower bounded by $L.\min_{i,j}S_{ij}$. With the appropriate $\psi$, $\psi(\mathbf{P})$ will be a low-rank matrix and hence Algorithm 3 and the associated recovery guarantee of Theorem 1 holds.
7 Experiments
In this section, we answer Question 5. We now perform simulations and experiments with LLM in order to understand the performance of our robust ranking approach in practice in both general and LLM preference dataset settings.
7.1 Evaluation Criterion
We use several evaluations to assess our proposed methods’ effectiveness against unobserved and adversarial corrupted comparisons.
7.1.1 Normalized Frobenius Error
First, To measure the relative error between two preference matrices in terms of their elements’ magnitudes, we use the normalized Frobenius error (NFE). NFE between two matrices $P$ and $\overline{P}$ is defined as:
$NFE(P,\overline{P})=\frac{\|P-\overline{P}\|_{Fro}}{\|P\|_{Fro}},$
where the Frobenius norm, denoted as $\|A\|_{Fro}$, for a matrix $A$ is calculated by:
$\|A\|_{Fro}=\sqrt{\sum_{i=1}^{n}\sum_{j=1}^{n}|a_{ij}|^{2}}$
In this formula, $a_{ij}$ represents the element of the matrix $A$ in the $i$th row and $j$th column. The Frobenius norm is the square root of the sum of the absolute squares of all elements in the matrix. Thus, the numerator $\|P-\overline{P}\|_{Fro}$ calculates the Frobenius norm of the difference between the original and reconstructed matrices, and the denominator $\|P\|_{Fro}$ calculates the Frobenius norm of the original matrix. The ratio provides a measure of the relative error normalized by the magnitude of the original matrix.
7.1.2 Correlation Coefficient
Second, we compute the correlation coefficient for corresponding elements in these matrices to assess the similarity between the original matrix $P$ and the reconstructed matrix $\overline{P}$. The correlation coefficient, denoted as $corr$, between the elements of these two matrices can be defined as:
$corr(P,\overline{P})=\frac{\sum_{i=1}^{n}(P_{i}-\langle P\rangle)(\bar{P}_{i}-\langle\bar{P}\rangle)}{\sqrt{\sum_{i=1}^{n}(P_{i}-\langle P\rangle)^{2}}\sqrt{\sum_{i=1}^{n}(\bar{P}_{i}-\langle\bar{P}\rangle)^{2}}},$
where $\langle P\rangle$ and $\langle\bar{P}\rangle$ denote the mean values of the elements within the $P$ and $\bar{P}$ matrices, respectively. $n$ represents the total number of elements in each matrix.
This formula quantifies the linear relationship between the matrices’ elements. A correlation coefficient close to $1$ indicates a strong positive linear relationship, whereas a value close to $-1$ suggests a strong negative linear relationship. A coefficient around 00 implies no linear relationship.
7.1.3 Ranking Distance
Third, for ease of reference, we rewrite the $\operatorname{dist}\left(\widehat{\sigma},\mathbf{P}\right)$ formula, which evaluates the distance between rankings obtained by corrupted and recovered matrices, previously defined in Section 3.1:
$\operatorname{dist}\left(\widehat{\sigma},\mathbf{P}\right)$ $:=\begin{pmatrix}n\\ 2\end{pmatrix}^{-1}\sum_{i<j}\mathbbm{1}\left((P_{ij}>1/2)\wedge(\widehat{\sigma}(i)\succ\widehat{\sigma}(j))\right)$
$+\begin{pmatrix}n\\ 2\end{pmatrix}^{-1}\sum_{i<j}\mathbbm{1}\left((P_{ji}>1/2)\wedge(\widehat{\sigma}(j)\succ\widehat{\sigma}(i))\right),$
where $\widehat{\sigma}$ is the global ranking after applying ranking procedure with $\bar{P}$.
7.2 Performance of Robust Ranking in General Setting
First, we begin with the BTL model. We generate synthetic pairwise comparison data and also adversarial sparse matrix as follows. We generate the entries of the BTL parameter vector $\mathbf{w}$ from $\mathcal{N}(0,\nu^{2})$ followed by generating the ground truth preference matrix from with $y_{ij}^{k}$ is sampled for all response pairs $(i,j)$ for a fixed $K$. The adversarial sparse matrix $\mathbf{S}$ is generated as a skew-symmetric matrix where each entry is non-zero independently with probability $d/n$ followed by generating a value for an entry from $U(5,10)$ and then setting the sign to be positive with probability $1/2$; this corruption matrix is then added to the $\psi(\mathbf{P})$ to give $\psi(\mathbf{P^{c}})$ which is then input to our algorithm; the same $\mathbf{P^{c}}$ is used for the other algorithms as well.
Figure 3: Robust recovery results of the BTL model: we fix $\nu=2$ and vary $d$ in the left plot; we fix $d=100$ and vary $\nu$ in the left plot. The black line represents RORATRON, the blue line represents Maximum Likelihood (ML), the pink line represents Rank Centrality (RC), and the red line represents Borda Count (BC).
We take the number of responses to be $n=500$. In plots in Figure 3, we compare the performance of our RORATRON approach using Algorithm 3 against well-known ranking algorithms, such as Rank Centrality (RC) (Negahban et al.,, 2012), Maximum Likelihood (ML) and Borda Count (BC) (Jiang et al.,, 2011), with special attention to robustness to the noise model that we consider in this paper. We vary two parameters, namely, $\nu$, the spread of the BTL scores, and $d$, the density of the adversarial corruption matrix. All our results averaged over five runs. We observe that our algorithm maintains low recovery error in spite of increasing the problem hardness, thus outperforming previous approaches in all cases.
7.3 Performance of Robust Ranking in LLM Preference Dataset
In this illustrative experiment, from the MT-Bench dataset (Zheng et al.,, 2023), we collect the data of the first prompt “Compose an engaging travel blog post about a recent trip to Hawaii, highlighting cultural experiences and must-see attractions" and its six responses from GPT-3.5, GPT-4 (OpenAI et al.,, 2023), Claude-v1 (Anthropic,, 2023), Vicuna-13B (Chiang et al.,, 2023), Alpaca-13B (Taori et al.,, 2023), and LLaMA-13B (Touvron et al., 2023a, ). Additionally, we generated nine responses to the same prompt using Llama-2-70B-chat-hf (Touvron et al., 2023b, ), Falcon-180B-chat Almazrouei et al., (2023), Openchat-3.5 (Wang et al.,, 2023), Mixtral-8x7B-Instruct-v0.1 (Jiang et al.,, 2024), Mistral-7B-Instruct-v0.2 (Jiang et al., 2023a, ), Gemini-pro (Gemini et al.,, 2023), Dolphin-2.2.1-mistral-7B (Hartford,, 2023), Solar-10.7B-instruct-v1.0 (Kim et al.,, 2023), Yi-34B-chat (01.ai,, 2023) from Hugging Face’s HuggingChat (Hugging Face,, 2023) and LMSYS’s Chatbot Arena (Zheng et al.,, 2023). So we have $n=15$ responses.
Next, we rank the responses using OpenAI’s GPT-4 Turbo GPT-4-1106-preview (OpenAI et al.,, 2023). This ranking helps us create the BTL parameter vector $\mathbf{w}$. We then sort this vector descendingly for visually accessible when building the corresponding preference matrix $\mathbf{P}\in\mathbb{R}^{n\times n}$. With (${n\choose 2}$ comparisons in $\mathbf{P}$, we randomly remove entries based on a specified deletion probability parameter, $dp$, to simulate unobserved comparisons. We then create an adversarial skew-symmetric sparse matrix, $\mathbf{S}$, using the given matrix $\mathbf{P}$ and an adversarial corruption probability parameter $ap$. When corruption is applied, it involves randomly selecting a value from $U(-5,5)$ and then adding to the $\mathbf{P}$ to give $\mathbf{P}^{c}$ which is then become the input of our algorithm. It’s important to note that $\mathbf{P}$ is a skew-symmetric matrix, any corruption must be applied to both $ij$ and $ji$ values.
Our experiment results visualized in Figure 4 show that $dp=10\%$ and $da=10\%$ can significantly affect the ranking of different models and the rank of the matrix when performing logit link transformation. The ranking can get altered quite badly when compared to the original matrix. Also, the logit link transformation of the corrupted matrix is high-rank which indicates that there are noises in the matrix. By using CURATRON to impute the missing comparisons and filter out the noisy sparse matrix, we successfully reconstruct the original matrix, which is low-rank when in logit link transformed form. As a result, we obtain the correct ranking. We also obtain noisy comparisons that can be used to identify responders with malicious intent and prevent them from continuing to alter results.
We now examine how our algorithm performs across different levels of unobserved and adversarially corrupted comparisons. In the plots shown in Figure 5, we compare the performance of our approach by varying two parameters, $dp$ and $ap$. We use NFE, correlation, and ranking distance as defined earlier in section 7.1. Our results are averaged over 5 runs. When there is no adversarial noise, we can recover the original $\mathbf{P}$ with no NFE and perfect correlation and ranking, even if $50\%$ of the comparison data was missing. This suggests that we may not need to collect all comparisons from humans to obtain the entire data. We observe that, with $n=15$, we only need to obtain about $50-55\%$ of the 105 comparisons and fill in the rest with our algorithm to achieve a strict $0\%$ NFE, perfect correlation, and ranking. On the other hand, when missing data is absent, our algorithm performs well with NFE of approximately $6\%$, even when $35\%$ of the comparison data is adversarially corrupted. When both adversarial noise and missing data are present, we can achieve a low NFE of around $4\%$ when both $15\%$ of the comparison data is missing and $15\%$ of adversarially corrupted comparisons (30% in total) affect $\mathbf{P}$.
Figure 4: Left: Original matrix. Middle: corrupted matrix. Right: reconstructed matrix. The corrupted matrix has 10% adversarial corruptions and 10% of unobserved comparisons. We use our CURATRON algorithm to successfully recover the original matrix.
Figure 5: Average over 5 runs of reconstruction error, correlation, and distance between reconstructed ranking and original ranking for different percentages of unobserved and adversarial comparisons.
7.4 Preference Dataset Recovery and its effect on LLM Performance
Our goal in this experiment is to assess the effectiveness of our proposed algorithms in reducing corruption in LLM Preference datasets. To achieve this, we design our experiment in the following manner. We utilize the pre-generated responses from 34 models that prioritize the helpfulness objective as the preference dataset available on MT-Bench. We transform this dataset into a pairwise comparison format, where we discard tied comparisons. This transformed dataset is called the original set. We then intentionally corrupt 20% of the data from the original set, which results in a corrupted set. We then apply our algorithm to this corrupted set, which results in a recovered set. Finally, we fine-tune three models from the pretrained OpenHermes-2.5-Mistral-7B (Teknium1,, 2023) with these three datasets using DPO.
We use the AlpacaEval benchmark (Li et al.,, 2023) to evaluate our models, which prioritizes measuring helpfulness. Our DPO-tuned LLM models answer 805 questions on various topics based on the AlpacaEval instructions. To evaluate our models, we compare their responses to the baseline responses generated by a GPT-4 Turbo model, using a referee GPT-4 Turbo model (OpenAI et al.,, 2023). We use the discrete non-loss rate as our evaluation metric, which is calculated using the formula:
[…]
Table 1: Comparison of DPO finetuned models on three datasets: baseline, corrupted, and recovered MT-Bench datasets (Zheng et al.,, 2023). Higher values are better for the overlap with the original set and non-loss rate; whereas, for training loss, lower values are better. Being closer to the original dataset size is better in terms of dataset size.
Model OpenHermes-2.5-Mistral-7B (Teknium1,, 2023) DPO Finetuning Dataset type Dataset size Overlap with the original set Average training loss AlpacaEval non-loss rate Original (baseline) set 31376 100% 0.4600 9.81% Corrupted set 44877 68.22% 0.5592 9.57% Recovered set 32391 96.65% 0.4638 10.06%
After comparing the data overlap of three datasets, we found that the recovered dataset has a 96.65% overlap with the original dataset, while the corrupted dataset only overlaps about 68.22% with the original datasets. Our algorithm helped us reduce the size of the recovered dataset by 27.8% (from 44,877 comparisons to 32,391 comparisons), which suggests that our algorithm is sufficient in recovering the original dataset (31,376 comparisons). We also observed that the training losses of the original and recovered datasets were about 10% lower on average than the corrupted dataset. Based on these insights, we expect similar performance in finetuning using the original and recovered datasets. Notably, the corrupted model achieved a non-loss rate of 9.57% against GPT-4 Turbo (OpenAI et al.,, 2023), while the original model achieved 9.81% and the recovered model achieved 10.06%.
8 Conclusion
Our study examines how missing information and distorted feedback can impact LLMs, potentially compromising their performance in terms of alignment with human values. We have proposed a robust algorithm for provably correct and efficient ranking responses in the BTL, LR, and general binary choice models. This robust ranking data is then input in the PL step. Further, we also handled the partially observed setting, wherein only some response pairs are compared, by integrating matrix completion techniques into our robust learning algorithm. In all cases, we provided statistical and computational guarantees using novel techniques. Through our comprehensive analysis, we hope to contribute to the ongoing discussion on AI safety by helping to create and scale LLMs/AGI models that align with human values and expectations. Some future research directions include tightening the recovery results for partially observed settings under weaker conditions (possibly using noisy-case extensions of (Yi et al.,, 2016)), exploring other notions of adversarial noise, and understanding the minimax optimal rates for ranking estimators under various noise models. We also plan to study the parametric non-active pairwise ranking setting, studying lower bounds and practical algorithms in the active setting similar to (Heckel et al.,, 2016). Furthermore, it would be interesting to investigate whether we can extend this approach to solve the entity corruption problem in retrieval models, as shown in (Naresh et al.,, 2022). Another research direction could be defining an alignment framework that expands DPO to various objective functions based on Rank Centrality (Negahban et al.,, 2017). Finally, we aim to examine the relationship between robust PL and model capacity, as this can shed light on the trade-offs between model complexity and generalization performance.
Appendix A Prompt template and answer of LLM judge
We evaluate various responses from different Language Models (LLMs) using GPT-4-1106-preview (OpenAI et al.,, 2023) as the judge. We have adapted the prompt template from MT-Bench (Zheng et al.,, 2023), but instead of scoring each candidate separately, we score all models in a single turn. Please note that the results generated by GPT-4-1106-preview judge may not be the same when the model gets updated. To ensure that the judge’s response is deterministic, we have set the temperature parameter to 0.
System prompt: You are a helpful assistant. Prompt Template: [Instruction] Please act as an impartial judge and evaluate the qualities of the responses provided by many AI assistants to the user question displayed below. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Begin your evaluation by providing a short explanation to each of the assistant responses. Be as objective as possible. Your judgment should be based solely on the content of the answers, disregarding factors like the position of the responses, the length of the responses, or the names of the assistants. After providing your explanation, you must rate the responses on a scale of 0 to 1 by strictly following this format: Final verdict for each assistant (full list): {‘Assistant A’: $score_{a}$, ’Assistant n’: $score_{n}$} For example: Final verdict for each assistant (full list): {‘Assistant A’: 0.30, ‘Assistant B’: 0.19, ‘Assistant C’: 0.75} [The Start of Assistant A’s Answer] Assistant A’s Answer [The End of A’s Answer] [The Start of Assistant n’s Answer] Assistant n’s Answer [The End of n’s Answer] Judge’s Answer: Final verdict for each assistant (full list):
{‘Assistant A’: 0.90, ‘Assistant B’: 0.95, ‘Assistant C’: 0.80, ‘Assistant D’: 0.75, ‘Assistant E’: 0.85, ‘Assistant F’: 0.70, ‘Assistant G’: 0.60, ‘Assistant H’: 0.65, ‘Assistant I’: 0.55, ‘Assistant J’: 0.40, ‘Assistant K’: 0.45, ‘Assistant L’: 0.50, ‘Assistant M’: 0.20, ‘Assistant N’: 0.35, ‘Assistant O’: 0.30}
Explanation of Scores: A scored well as it provides a structured, detailed, and vivid account of the Hawaiian experience, including cultural and natural attractions. B earned the highest score for a well-written, engaging narrative combining in-depth cultural experiences with comprehensive overviews of natural attractions, presenting it all in an engaging manner. C’s response was immersive, providing a daily itinerary format which gives readers a chronological perspective but loses some detail and depth in the descriptions, resulting in a slightly lower score. D’s response was engaging, offering personal reflections and detailed descriptions of the experiences, but it could capitalize on a more informative angle regarding cultural aspects. E crafted an elaborate and expressive narrative with a strong focus on cultural elements and natural beauty. It’s highly descriptive, which enhances its appeal. F offered a nice overview with personal insights but was more generic compared to others, with less distinctive detail on the cultural elements and attractions. G provided a good variety of information but lacked a personal touch and the engaging storytelling present in others, hence a middle-of-the-road score. H put forth a decent attempt to cover cultural experiences and attractions, but the content was less rounded and not as rich in personal perspective. I’s storytelling lacked in-depth detail and read more like an overview, resulting in a lower rating compared to others that provided richer narratives. J’s response was brief and lacked the depth, detail, and storytelling elements found in the higher-scoring responses. K focused more on personal experiences rather than providing informative cultural insights, which resulted in a score below others that balanced both elements better. L offered an informative narrative but was less dynamic and less vivid in description than higher-scoring answers. M’s response was too factual and lacked the storytelling and descriptive engagement present in other responses, leading to a lower score. N provided a decent overview but lacked the richness of storytelling and description, leaving it towards the lower end of the scale. O’s draft was informative but lacked the detail and narrative quality that made other responses stand out, resulting in a lower score.